OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. - tan2x + sec2x = 1

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Heyy!

OpenStudy (misty1212):

is that \[-\tan^2(x)+\sec^2(x)=1\]?

OpenStudy (anonymous):

Yes.

OpenStudy (misty1212):

ok lets start with the mother of all trig identities \[\sin^2(x)+\cos^2(x)=1\]

OpenStudy (misty1212):

divide each term by \(\cos^2(x)\)

OpenStudy (misty1212):

you get \[\frac{\sin^2(x)}{\cos^2(x)}+\frac{\cos^2(x)}{\cos^2(x)}=\frac{1}{\cos^2(x)}\\ \tan^2(x)+1=\sec^2(x)\]

OpenStudy (misty1212):

subtract \(\tan^2(x)\) from both sides and you get exactly what you want \[-\tan^2(x)+\sec^2(x)=1\]

OpenStudy (unknownunknown):

I like your smile Misty.

OpenStudy (anonymous):

i'd walk a mile...

OpenStudy (misty1212):

you good with that?

OpenStudy (misty1212):

@unknownunknown thanks! i like @clairvoyant1 smile too

OpenStudy (unknownunknown):

Yes, you both have a beautiful smile.

OpenStudy (anonymous):

Yes, so 1 - sin^2(x) = cos^2(x)

OpenStudy (misty1212):

yeah you can use that too then divide by \(\cos^2(x)\) to get what you want

OpenStudy (anonymous):

-sin^2(x)/cos^2(x) + 1/cos^2(x) ?

OpenStudy (misty1212):

you got \[1-\sin^2(x)=\cos^2(x)\] right ?

OpenStudy (misty1212):

oh yeah, what you wrote except you forgot the right hand side of the equal sign

OpenStudy (anonymous):

yes.

OpenStudy (misty1212):

ok so the first term you wrote above is \(-\frac{\sin^2(x)}{\cos^2(x)}\) that is the same as \(-\tan^2(x)\)

OpenStudy (misty1212):

the second term you wrote is \(\frac{1}{\cos^2(x)}\) which is \(\sec^2(x)\)

OpenStudy (misty1212):

and on the right side of the equal sign you would have \(\frac{\cos^2(x)}{\cos^2(x)}\) which is \(1\)

OpenStudy (anonymous):

@clairvoyant1 you got that? misty wrote it all out should be good right?

OpenStudy (anonymous):

and that's the rhs right?

OpenStudy (anonymous):

yes the rhs is 1

OpenStudy (anonymous):

so you are done at that step with \[-\tan^2(x)+\sec^2(x)=1\] which is what you wanted to show

OpenStudy (dan815):